Periodic Inventory Using FIFO, LIFO, and Weighted Average Cost Methods

The units of Manganese Plus available for sale during the year were as follows:

Mar. 1 Inventory 22 units @ $29 $638
June 16 Purchase 28 units @ $33 924
Nov. 28 Purchase 43 units @ $38 1,634
93 units $3,196

There are 12 units of the product in the physical inventory at November 30. The periodic inventory system is used. Round answers to the nearest whole dollar.

a. Determine the inventory cost by the FIFO method.
$

b. Determine the inventory cost by the LIFO method.
$

c. Determine the inventory cost by the average cost methods.
$
Answer #1
a
Inventory cost by the FIFO method 456 =12*38
b
Inventory cost by the LIFO method 348 =12*29
c
Inventory cost by the average cost method 412 =3196*12/93